# Reversibility in finite-dimensional collapse dynamics

### v2

@@ Abstract
We study finite dimensional quantum systems with arbitrary collapse events, establishing, under no-information-erasure conditions, a structural no-go for trajectory-level operational irreversibility. More precisely, we fix a realization map (a physically admissible selector of the collapse dynamics) and do not rely on any regularity of the induced dynamics. We prove that, for every realization of the collapse dynamics, there exists a topologically closed, forward-invariant subset of the projective state space on which any two states can be connected with arbitrarily fine Fubini–Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness, explicit erasure, or coupling to reservoirs.
@@ 1.1 Reversibility in unitary versus collapse dynamics
In unitary quantum dynamics of finite-dimensional systems, recurrence phenomena, either in the measurable or topological sense, arise naturally from the continuity of unitary evolution. Indeed, since the Schrödinger dynamics is continuous (indeed isometric), in the compact case the classical recurrence theorems by Poincaré [poincare1890probleme, bocchieri1957quantum] and Birkhoff [birkoff1912quelques] ensure, respectively, that almost every orbit admits arbitrarily close returns in state space, and that quasi–periodic orbits exist. Therefore, within purely unitary evolution of compact systems, the dynamics admits no intrinsic arrow of time.
@@
When wavefunction collapse is included in the picture, this structure breaks. For the discrete–time setting, a natural analogue of the previously described unitary evolution is obtained by following a single, consistent branch of the dynamics in an open system, an approach that has been introduced for the first time (within quantum optics) in [dalibard1992wave], and subsequently elaborated in many other works (for instance [carmichael1993open] and [wiseman2009quantum], the latter also discussing single-branch dynamics from an epistemological point of view). Consider a finite-dimensional space $\mathbb{P}(\mathcal{H})$ of pure states, a single observable $A$ with finite spectrum, and a particular state $\psi$. At each time step either no collapse occurs (so one applies a unitary Schrödinger evolution $\psi\mapsto U\psi$) or a projective collapse $\{P_{i}\}$ occurs. The future outcome itinerary that is realized by $\psi$, that we indicate by $\omega=\{i_{0},i_{1},\dots\}$ ($i_{k}$ being symbolic tools representing the possible collapse outcomes as well as the “blank”, no-collapse case), has to be of course admissible from a physical point of view. In particular, each selected outcome must have a nonzero Born weight for the state immediately preceding that step. If we assign consistently a selected outcome to every state, the evolution along the outcome trajectory provides a map $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ that is, in general, many–to–one and highly discontinuous, because there is no justification in assuming that close states will evolve into close collapsed states. Arbitrarily close states are allowed, in general, to collapse to eigenstates belonging to distinct eigenspaces of the chosen observable; this is not at all surprising if one considers that even the same pre–collapse state may be sent to different post–collapse eigenstates at distinct collapse events. Moreover, the eigenspaces of the observable are not absorbing for $T$ unless $U$ preserves them (i.e. $U,P_{i}]=0$); between collapses, the unitary evolution typically carries states out of these subspaces. Consequently, the hypotheses behind the Poincaré/Birkhoff recurrence theorems are violated on the observed system. [^2] If one re–includes apparatus and environment as a closed finite-dimensional composite, the global pure state again evolves unitarily on a compact projective space, and recurrence can be recovered.
@@
In the classical compact setting (where the Poincaré–Birkhoff recurrence theorems apply), one studies the long–time behavior of a single map on a compact state space. We want to explore the closest possible analogue when generalizing unitary evolution of closed systems to collapse dynamics of open systems. We fix a selector $\omega$ of the skew-product dynamics and thus obtain a single branch on the projective state space. Notice that this choice cleanly separates the predictability issue (which outcome occurs) from the regularity issue (whether the induced map $T$ is continuous). Moreover, we take the selector in its most literal observational sense:
@@
    every outcome itinerary has to be physically admissible and time–compatible for each initial state;
@@
We remark that is not the standard stance in the quantum–trajectories literature: the seemingly intractable irregularity, that such state–wise selectors permit, is usually avoided by averaging over outcomes or retaining stochasticity and works with the skew product or with stochastic master equations (diffusive/jump SMEs), where continuity and stability are recovered in measure or in expectation rather than pointwise for a deterministic selector–induced map [carmichael1993open, belavkin1992quantum, barchielli2009quantum, attal2006repeated, wiseman2009quantum]. Our approach is instead to explore what can be said, from the point of view of recurrence, even allowing full irregularity. We fix thus a realization map (the chosen selector of the dynamics) assuming only the (physics-dictated) properties 1-2 above, and allowing the density property (3), which implies everywhere discontinuity for the induced evolution map $T$.
@@ 1.2 The result
Landauer’s principle (see [landauer1961irreversibility]) bounds the minimal heat cost of logically irreversible operations (such as erasure), and conversely permits, in principle, arbitrarily low dissipation for logically reversible transformations in the quasi–static limit.
The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox [bennett1982thermodynamics, bennett2003notes, hanson2018landauer]. In particular, we share the following position, expressed by Bennett (from [bennett1982thermodynamics]):
@@
Our realization map, indeed, does not erase (or coarse-grain) outcome information: the entire future outcome sequence is retained as a fixed record, and collapse consists solely in following one realized branch.
@@ 1.3 Where a naive diagonal grid argument fails
We implement (in Appendix) a transfinite argument that enforces both spatial nesting and dynamical coherence across scales: at successor stages it follows the dynamics induced by the selector; at limit stages it passes to accumulation points of cofinal sequences, gluing thus information from all previous scales. This procedure yields a strongly chain–recurrent point, and even a topologically closed TT–invariant subset that is internally $\mathcal{SC}_{d_{FS}}$–transitive. We remark that the procedure is constructive: we do not use any form of the Axiom of Choice and the two basic operations considered (taking images of the map and passing to an accumulation point of an orbit) clearly have finite-precision analogues.
@@ 2 Preliminaries
2.1 Notation and definitions for Quantum Dynamics formalism
@@
We recall the definition of the Fubini-Study metric (for a more detailed discussion, see [bengtsson2017geometry]).
@@
2.2 Notation and definitions for abstract Topological Dynamics formalism
In this Section we provide the topological dynamical concepts that will be used to prove our main result. Most of them appear only in the proof of Theorem 4.5, given in the Appendix.
We say that $(X,f)$ is a topological dynamical system if $X$ is a topological space with a metric $d:X^{2}\to\mathbb{R}_{0}^{+}$ compatible with the topology on $X$ and $f:X\to X$ a map. By $f^{n}(x)$ with $x\in X$ and $n\in\mathbb{N}$ we mean the $n$-times composition $f\circ f\circ\ldots\circ f(x)$. When we say that $(X,f)$ is a compact dynamical systemif we require, in addition, that $X$ is a compact metric space.
Let us recall the notion of $\epsilon$-chain or $\epsilon$-pseudo-orbit [kurka2003topological, p. 48].
@@ 2.2 Notation and definitions for abstract Topological Dynamics formalism
Recall the concept of strong chain-recurrence originally introduced by Easton [easton2006chain].
@@
3 Discrete-time Quantum Dynamics
@@
Note that we are not assuming (3), but we stress that we are allowing it. Stronger formulations (for instance, requiring a positive Fubini–Study measure or open support of each outcome set) would impose quantitative rather than merely topological versions of the same idea, but note that, even under this minimal topological condition, the map $T_{U,D}$ is discontinuous everywhere.
@@
Note that the existence of a realization map does not conflict with the Kochen–Specker (KS) no-go theorem [kochen2011problem]. In fact, it addresses a completely different point. KS forbids noncontextual, dispersion-free value assignments to all projections simultaneously (in dimension $\geq 3$), preserving functional relations across incompatible observables. By contrast, throughout we fix a single measurement context $\{P_{j}\}_{j\in\Lambda}$ (one PVM), and the realization map $D:\mathbb{P}(\mathcal{H})\to\Omega$ assigns an outcome sequence only for this context. The assignment is explicitly contextual/history-dependent and makes no claims about outcomes for other, incompatible PVMs. Hence $D$ is not a global valuation on the projection lattice, and the hypotheses of KS are not met.
@@ 4 Reversibility
The rank-2 operator $K(v,\delta)$ is clearly skew-Hermitian and it is a rotational generator of angle $\delta$ on the plane, i.e. it verifies $e^{K(v,\delta)}w=v.$ Moreover, $\|K(v,\delta)\|=\delta=O(\epsilon)$. Let us now set the following
@@
The following result is key to our reasoning. For the proof, see Appendix. The proof is a refinement of that of Theorem 3.9 in [della2025chains]. To make the present work independent, we cover it in detail without relying on the proof of the weaker result.
@@
By Theorem 4.5, a sufficient condition for having an infinite chain-transitive subsystem is that $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ has no periodic orbits, as no finite set of states can be chain-transitive unless there are periodic cycles.
@@
This purely topological argument is, of course, far from optimal. The existence of periodic orbits does not prevent at all the existence of an infinite chain-transitive subset, and the rough estimate by cardinality can be improved. The question of how large is the largest chain-transitive subsystem is explored in the forthcoming work [FLS_].
@@ 5 Conclusions
Our work establishes a structural no-go for operational irreversibility along individual realized branches, under finite-dimensional, no-erasure conditions. Specifically, Theorems 4.5 and 4.6 tell us the following fact (which was indicated in the Introduction as our main result): in every finite-dimensional, open quantum system, there is a topologically closed, invariant (under the realization-induced dynamics) subsystem $S$ where any two states are connected via pseudo-orbits with arbitrarily small Fubini-Study error and energetic cost. By Def. 1.1, this means that $S$ is operationally reversible: no operational arrow of time arises.
@@
It is perhaps useful to frame this result by analogy with the logical structure of some classical thermodynamics concepts. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law, including information in the entropy balance, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands, structurally, in a similar relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (finite dimensionality and compactness) under which the bound is asymptotically saturated. Observe, however, that Carnot’s theorem is an inherently stronger result, within classical thermodynamics, than ours within quantum reversibility theory. Indeed, Carnot’s theorem provides a universal criterion, whereas our result only identifies a suitable subsystem showing quasi-reversibility.
@@
Let us finally remark that our argument crucially relies on the retention of the infinite record of outcomes provided by the realization map. This infinite history is essential because, to steer the system toward a target state with arbitrarily small FS error $\epsilon>0$, the perturbations must be tailored to the specific sequence of realized collapse outcomes encountered along the chain. Although only finitely many corrections are required to achieve any fixed $\epsilon$-precision, the number and nature of these corrections generally become unbounded as $\epsilon\to 0$, which means that, in principle, one needs the access to the entire future itinerary to prescribe the appropriate steering strategy at every scale. A similar observation is key in Bennett’s resolution of Maxwell’s demon paradox, where he showed that a demon’s finite memory forces logically irreversible erasure to accommodate new data, thereby incurring an entropic cost that preserves the second law [bennett1982thermodynamics]. Our no-erasure framework mirrors the idealized demon scenario, where an infinite memory sidesteps dissipation. Yet, as Bennett argued for any physical demon, real-world implementations face finite-size constraints that inevitably lead to erasure. Analogously, any practical realization of our procedure with a truncated record would introduce operational irreversibility, which recovers the second law.
@@
In this Appendix we address the proof of Theorem 4.5.
We start recalling the following result that shows that in a compact dynamical system, without any regularity assumption on the map, there always exists a generalized recurrent point for the system, that is, a point $x\in X$ such that $x\in\bigcap_{d\in M}\mathcal{SCR}_{d}(f)=:GR(f)$ where $M$ is the set of the metrics compatible with the topology of $X$.
@@
[della2025chains, Theorem 3.5] Let $(X,f)$ be a compact dynamical system. Then $GR(f)\neq\emptyset$.
@@
Assume that every point in the space $X$ has an infinite orbit, otherwise the claim of Theorem 5.1 follows easily. Pick $x_{0}:=x\in X$ and set $x_{k}:=f(x_{k-1})$ for all $k\in\mathbb{N}$. Take $y\in\mathcal{O}^{\prime}(x)$, which is nonempty due to the compactness of $X$, and set $x_{\omega}:=y$. Let $\alpha>\omega$ be an ordinal number and assume that $x_{\beta}$ has been defined for every $\beta<\alpha$. Define the following sets:
@@
and $x_{\alpha}:=y$ with $y$ any point in $S^{*}_{\alpha}(x)$. Notice that $S^{*}_{\alpha}(x)\neq\emptyset$ (see, e.g. [munkres2013topology, Th. 26.9, p. 169]), since $|S_{\beta,\alpha}(x)|\geq\aleph_{0}$ for every $\beta<\alpha$ so that $(S_{\beta,\alpha}(x))^{\prime}$ is a closed and nonempty set. The second step in the proof of [della2025chains, Theorem 3.5] is to show that for every ordinal $\alpha$,
@@
Moreover, by exploiting the proof of property (9) in [della2025chains] we have the following
@@
For every $\beta<\eta\leq\alpha$ and every $\epsilon>0$, there is a strong $(\epsilon,d)$-chain from $x_{\beta}$ to $x_{\eta}$ whose points belong to $S_{\beta,\eta+1}(x)$. Therefore, if $x_{\alpha}=x_{\beta}$ for some $\alpha>\beta$, then by the transitivity of $\mathcal{SC}_{d}$ we have that the set $S_{\beta,\alpha}(x)$ is internally $\mathcal{SC}_{d}$-transitive, that is, for every $y,z\in S_{\beta,\alpha}(x)$ and for every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $y$ to $z$ whose points belong to $S_{\beta,\alpha}(x)$.
The following lemma is a direct consequence of [della2025chains, Lemma 3.6 and Remark 3.8] and Remark 5.2.
@@
In particular, by Lemma 5.1, we can say that for $x\in X$, there exists a least ordinal $\alpha>0$ such that $x_{\alpha}\in S_{\alpha}(x)$ and so, in particular, $x_{\alpha}=x_{\beta}$ for some $\beta<\alpha$. Moreover, the set $S_{\beta,\alpha}(x)$ is an invariant and internally $\mathcal{SC}_{d}$-transitive subset of $X$.
In [della2025chains] it is shown that a compact dynamical system has a chain-transitive subsystem. Using the previous lemma and Remark 5.2, we want to prove that every compact dynamical system has, in particular, a $\mathcal{SC}_{d}$-transitive subsystem.
Now are ready to prove the result, that we formulate here in the abstract topological dynamical notation.
@@
We need to show that there is a closed, invariant and internally $\mathcal{SC}_{d}$-transitive subset of $X$. Suppose, towards a contradiction, that $(X,f)$ does not have a $\mathcal{SC}_{d}$-transitive subsystem. In particular, we can assume that every orbit is infinite. Let $x^{0}\in X$. By Lemma 5.1, there exists a least ordinal $\alpha_{0}>0$ such that $x^{0}_{\alpha_{0}}\in S_{\alpha_{0}}(x^{0})$, and so $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$. By Lemma 5.3 and Remark 5.2, the set $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset. Since $S_{\beta_{0},\alpha_{0}}(x^{0})$ cannot be closed, there exists some $y\in(S_{\beta_{0},\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\beta_{0},\alpha_{0}}(x^{0})$. Let $x\in S_{\alpha_{0}}(x^{0})$ and $\epsilon>0$, we want to show that there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$. Let $\beta_{0}\leq\eta<\alpha_{0}$ be such that $d(x_{\eta},y)<\epsilon/2$. By property (9) and Remark 5.2, it follows that there exists a strong $(\epsilon/2,d)$-chain $x=y_{0},y_{1},\ldots,y_{n-1},y_{n}=x_{\eta}$ from $x$ to $x_{\eta}$ whose points belong to $S_{\alpha_{0}}(x^{0})$. By the triangle inequality,

### v3

@@ Abstract
We study finite dimensional quantum systems with arbitrary collapse events, establishing, under no-information-erasure conditions, a structural no-go for trajectory-level operational irreversibility. More precisely, we prove that, for every physically admissible selector of the collapse dynamics, there exists a topologically closed, forward-invariant subset of the projective state space on which any two states can be connected with arbitrarily fine Fubini–Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch of outcomes guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness or information erasure.
@@ 1.1 Reversibility in unitary versus collapse dynamics
In unitary quantum dynamics of finite-dimensional systems, recurrence phenomena, either in the measurable or topological sense, arise from the continuity of unitary evolution. Indeed, since the Schrödinger dynamics is continuous (indeed isometric), in the compact case the classical recurrence theorems by Poincaré [20, 8] and Birkhoff [7] ensure, respectively, that almost every orbit admits arbitrarily close returns in state space, and that quasi–periodic orbits exist. Therefore, within purely unitary evolution of compact systems, the dynamics admits no intrinsic arrow of time.
@@
When wavefunction collapse is included in the picture, this structure breaks. For the discrete–time setting, a natural analogue of the previously described unitary evolution is obtained by following a single, consistent branch of the dynamics in an open system, an approach that has been introduced for the first time (within quantum optics) in [10], and subsequently elaborated in many other works (for instance [9] and [21], the latter also discussing single-branch dynamics from an epistemological point of view). Consider a finite-dimensional space $\mathbb{P}(\mathcal{H})$ of pure states, a single observable $A$ with finite spectrum, and a particular state $\psi$. At each time step either no collapse occurs (so one applies a unitary Schrödinger evolution $\psi\mapsto U\psi$) or a projective collapse $\{P_{i}\}$ occurs. The future outcome itinerary that is realized by $\psi$, that we indicate by $\omega=\{i_{0},i_{1},\dots\}$ ($\{i_{k}\}_{k=0,\dots N}$ being symbolic tools representing the possible collapse outcomes as well as the “blank”, no-collapse case), has to be of course admissible from a physical point of view. In particular, each selected outcome must have a nonzero Born weight for the state immediately preceding that step. If we assign a selected outcome to every state, the evolution along the outcome trajectory provides a map $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ that is, in general, many–to–one and highly discontinuous, because there is no justification in assuming that close states will evolve into close collapsed states. Arbitrarily close states are allowed, in general, to collapse to eigenstates belonging to distinct eigenspaces of the chosen observable; this is not at all surprising if one considers that even the same pre–collapse state may be sent to different post–collapse eigenstates at distinct collapse events. Moreover, the eigenspaces of the observable are not absorbing for $T$ unless $U$ preserves them (i.e. $U,P_{i}]=0$); between collapses, the unitary evolution typically carries states out of these subspaces. Consequently, the hypotheses behind the Poincaré/Birkhoff recurrence theorems are violated on the observed system. [^2] Of course if one re–includes apparatus and environment as a closed finite-dimensional composite, the global pure state again evolves unitarily on a compact projective space, and recurrence can be recovered.
@@
In the classical compact setting (where the Poincaré–Birkhoff recurrence theorems apply), one studies the long–time behavior of a single map on a compact state space. We want to explore the closest possible analogue when generalizing unitary evolution of closed systems to collapse dynamics of open systems. We fix a selector $\omega$ of the skew-product dynamics and thus obtain a single branch on the projective state space. This choice cleanly separates the predictability issue (which outcome occurs) from the regularity issue (whether the induced map $T$ is continuous or has any regularity). Moreover, we take the selector in its most literal observational sense:
@@
    every outcome itinerary has to be physically admissible and the histories seen by different states must be time–compatible;
@@
We remark that is not the standard stance in the quantum–trajectories literature: the seemingly intractable irregularity, that such state–wise selectors permit, is usually avoided, for instance by averaging over outcomes or retaining stochasticity, where continuity and stability are recovered in measure or in expectation rather than pointwise for a deterministic selector–induced map [9, 3, 2, 1, 21]. Our approach is instead to explore what can be said, from the point of view of recurrence, assuming only 1. and 2. above, so in particular allowing 3. We fix thus a realization map (the chosen selector of the dynamics) assuming only the (physics-dictated) properties 1-2 above, and allowing the density property (3), which implies everywhere discontinuity for the induced evolution map $T$.
@@ 1.2 The result
Landauer’s principle (see [18]) bounds the minimal heat cost of logically irreversible operations (such as erasure), and conversely permits, in principle, arbitrarily low dissipation for logically reversible transformations in the quasi–static limit.
The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox [5, 6, 13]. In particular, we share the following position, expressed by Bennett (from [5]):
@@
Our realization map, indeed, does not erase (or coarse-grain) outcome information: the entire outcome sequence is always available as a fixed record, and collapse consists solely in following one realized branch.
@@ 1.3 Where a naive diagonal grid argument fails
At the opposite extreme of the naive attempt just described, one could obtain an abstract pure existence result on the existence of a chain-transitive subsystem by using Zorn’s lemma on topologically closed, invariant subsets ordered by inclusion. Our aim, however, is to provide a constructive procedure, which is the only one, in our opinion, that works well with the operational character of the employed notion of reversibility. Therefore, not relying on the Axiom of Choice, we implement (in the Appendix) a transfinite argument that enforces both spatial nesting and dynamical coherence across scales. This procedure yields a strongly chain–recurrent point, and even a topologically closed TT–invariant subset that is internally $\mathcal{SC}_{d_{FS}}$–transitive (see Sections 2 and 3 for the relevant definitions). We remark that the two basic operations considered in the procedure, that is taking images of the map and passing to an accumulation point of an orbit, clearly have finite-precision analogues.
The paper is organized as follows: in Section 2 we introduce basic concepts and notation for the unitary evolution of the system and the abstract topological dynamics concepts that will be used later; in Section 3 we describe the selector of the collapse dynamics and the induced discrete-time evolution; in Section 4 we study reversibility for arbitrary (physically admissible) realization maps; Section 5 draws some conclusions; the Appendix is devoted to the proof of Theorem 4.5.
@@ 2 Preliminaries
2.1 Notation and definitions for the unitary evolution
@@
We recall the definition of the Fubini-Study metric (for a more detailed discussion, see [4]).
@@
2.2 Notation and definitions for Topological Dynamics formalism
In this Section we provide the topological dynamical concepts that will be used to prove our main result. We say that $(X,f)$ is a topological dynamical system if $X$ is a topological space with a metric $d:X^{2}\to\mathbb{R}_{0}^{+}$ compatible with the topology on $X$ and $f:X\to X$ a map. By $f^{n}(x)$ with $x\in X$ and $n\in\mathbb{N}$ we mean the $n$-times composition $f\circ f\circ\ldots\circ f(x)$. When we say that $(X,f)$ is a compact dynamical systemif we require, in addition, that $X$ is a compact metric space.
Let us recall the notion of $\epsilon$-chain or $\epsilon$-pseudo-orbit [17, p. 48].
@@ 2.2 Notation and definitions for Topological Dynamics formalism
Recall the concept of strong chain-recurrence originally introduced by Easton [12].
@@
3 Discrete-time dynamics
@@
Note that we are not assuming (3), but we stress that we are allowing it. Note that, under this topological condition, the map $T_{U,D}$ is discontinuous everywhere.
@@
Note that the existence of a realization map does not conflict with the Kochen–Specker (KS) no-go theorem [15]. In fact, it addresses a completely different point. KS forbids noncontextual, dispersion-free value assignments to all projections simultaneously (in dimension $\geq 3$), preserving functional relations across incompatible observables. By contrast, throughout we fix a single measurement context $\{P_{j}\}_{j\in\Lambda}$ (one PVM), and the realization map $D:\mathbb{P}(\mathcal{H})\to\Omega$ assigns an outcome sequence only for this context. The assignment is explicitly contextual/history-dependent and makes no claims about outcomes for other, incompatible PVMs. Hence $D$ is not a global valuation on the projection lattice, and the hypotheses of KS are not met.
@@ 4 Reversibility
The rank-2 operator $K(v,\delta)$ is clearly skew-Hermitian and it is a rotational generator of angle $\delta$ on the plane, i.e. it verifies $e^{K(v,\delta)}w=v.$ Moreover, $\|K(v,\delta)\|=\delta=O(\epsilon)$. We set:
@@
The following result is key to our reasoning. For the proof, see the Appendix. The proof is a refinement of that of Theorem 3.9 in [11]. To make the present work independent, we cover it in detail without relying on the proof of the weaker result.
@@
By Theorem 4.5, a sufficient condition for having an infinite chain-transitive subsystem is that $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ has no periodic orbits, as no finite set of states can be chain-transitive unless there exist periodic cycles.
@@
This purely topological argument is, of course, far from optimal. The existence of periodic orbits does not prevent at all the existence of an infinite chain-transitive subset, and the rough estimate by cardinality can be improved. The question of how large is, generically, the largest chain-transitive subsystem is explored in a forthcoming work (Farotti, M., Lucamarini, C., Steele, T.: Chain-transitive subsets for irregular compact systems [in preparation]).
@@ 5 Conclusions
Our work establishes a no-go for operational irreversibility along individual realized branches, under finite-dimensional, no-erasure conditions. Specifically, Theorems 4.5 and 4.6 tell us the following fact (which was indicated in the Introduction as our main result): in every finite-dimensional, open quantum system, there is a topologically closed, invariant (under the realization-induced dynamics) subsystem $S$ where any two states are connected via pseudo-orbits with arbitrarily small Fubini-Study error and energetic cost. By Def. 1.1, this means that $S$ is operationally reversible: no operational arrow of time arises.
@@
It is perhaps useful to frame this result by analogy with the logical structure of some classical thermodynamics concepts. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law, including information in the entropy balance, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands in a similar relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (compactness) under which the bound is asymptotically saturated. Observe, however, that Carnot’s theorem is an inherently stronger result, within classical thermodynamics, than ours within quantum reversibility theory. Indeed, Carnot’s theorem provides a universal criterion, whereas our result only identifies a suitable subsystem showing quasi-reversibility.
@@
Let us finally remark that our argument crucially relies on the retention of the infinite record of outcomes provided by the realization map. This infinite history is essential because, to steer the system toward a target state with arbitrarily small FS error $\epsilon>0$, the perturbations must be tailored to the specific sequence of realized collapse outcomes encountered along the chain. Although only finitely many corrections are required to achieve any fixed $\epsilon$-precision, the number of these corrections becomes unbounded, in general, as $\epsilon\to 0$, which means that, in principle, one needs the access to the entire future itinerary to prescribe the appropriate steering strategy at every scale. A similar observation is key in Bennett’s resolution of Maxwell’s demon paradox, where he showed that a demon’s finite memory forces logically irreversible erasure to accommodate new data, thereby incurring an entropic cost that preserves the second law [5]. Our no-erasure framework mirrors the idealized demon scenario, where an infinite memory sidesteps dissipation. Yet, as Bennett argued for any physical demon, real-world implementations face finite-size constraints that inevitably lead to erasure. Analogously, any practical realization of our procedure with a truncated record would introduce operational irreversibility, which recovers the second law.
Declarations
The authors declare that they have no conflicts of interest.
The manuscript has no associated data.
@@
In this Appendix, we address the proof of Theorem 4.5. As stated in the Introduction, we aim to provide an argument that does not rely on the Axiom of Choice, so that the construction makes the link between the chain-transitive subsystem and the dynamics transparent.
We start recalling the following result that shows that in a compact dynamical system, without any regularity assumption on the map, there always exists a generalized recurrent point for the system, that is, a point $x\in X$ such that $x\in GR(f):=\bigcap_{d\in M}\mathcal{SCR}_{d}(f)$ where $M$ is the set of the metrics compatible with the topology of $X$.
@@
[11, Theorem 3.5] Let $(X,f)$ be a compact dynamical system. Then $GR(f)\neq\emptyset$.
@@
Assume that every point in the space $X$ has an infinite orbit; otherwise, the claim of Lemma 5.1 follows easily. Pick $x_{0}:=x\in X$ and set $x_{k}:=f(x_{k-1})$ for all $k\in\mathbb{N}$. Take $y\in\mathcal{O}^{\prime}(x)$, which is nonempty due to the compactness of $X$, and set $x_{\omega}:=y$. Let $\alpha>\omega$ be an ordinal number and assume that $x_{\beta}$ has been defined for every $\beta<\alpha$. Define the following sets:
@@
and $x_{\alpha}:=y$ with $y$ any point in $S^{*}_{\alpha}(x)$. Notice that $S^{*}_{\alpha}(x)\neq\emptyset$ (see, e.g. [19, Th. 26.9, p. 169]), since $|S_{\beta,\alpha}(x)|\geq\aleph_{0}$ for every $\beta<\alpha$ so that $(S_{\beta,\alpha}(x))^{\prime}$ is a closed and nonempty set. The second step in the proof of [11, Theorem 3.5] is to show that for every ordinal $\alpha$,
@@
Moreover, by exploiting the proof of property (9) in [11] we have the following
@@
For every $\beta<\eta\leq\alpha$ and every $\epsilon>0$, there is a strong $(\epsilon,d)$-chain from $x_{\beta}$ to $x_{\eta}$ whose points belong to $S_{\beta,\eta+1}(x)$. Hence, if $x_{\alpha}=x_{\beta}$ for some $\alpha>\beta$, then by the transitivity of $\mathcal{SC}_{d}$ we have that the set $S_{\beta,\alpha}(x)$ is internally $\mathcal{SC}_{d}$-transitive, that is, for every $y,z\in S_{\beta,\alpha}(x)$ and for every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $y$ to $z$ whose points belong to $S_{\beta,\alpha}(x)$.
The following lemma is a direct consequence of [11, Lemma 3.6 and Remark 3.8] and Remark 5.2.
@@
The proof of this lemma is a consequence of the construction in Lemma 5.1, and by using [11, Remark 3.8] and Remark 5.2, it is analogous to the proof of [11, Lemma 3.6]. In particular, from Lemma 5.3, we can say that for every $x\in X$, there exists a least ordinal $\alpha>0$ such that $x_{\alpha}\in S_{\alpha}(x)$ and thus $x_{\alpha}=x_{\beta}$ for some $\beta<\alpha$. Moreover, the set $S_{\beta,\alpha}(x)$ is an invariant and internally $\mathcal{SC}_{d}$-transitive subset of $X$.
In [11], it is shown that a compact dynamical system has a chain-transitive subsystem. Using Lemma 5.3 and Remark 5.2, we want to prove that every compact dynamical system has, in particular, a $\mathcal{SC}_{d}$-transitive subsystem.
Now we are ready to prove Theorem 4.5, that for consistency we state here in abstract topological dynamical terms.
@@
We need to show that there is a closed, invariant, and internally $\mathcal{SC}_{d}$-transitive subset of $X$. Suppose, towards a contradiction, that $(X,f)$ does not have a $\mathcal{SC}_{d}$-transitive subsystem. In particular, we can assume that every orbit is infinite. Let $x^{0}\in X$. By Lemma 5.1, there exists a least ordinal $\alpha_{0}>0$ such that $x^{0}_{\alpha_{0}}\in S_{\alpha_{0}}(x^{0})$, and so $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$. By Lemma 5.3 and Remark 5.2, the set $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset. Since $S_{\beta_{0},\alpha_{0}}(x^{0})$ cannot be closed, there exists some $y\in(S_{\beta_{0},\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\beta_{0},\alpha_{0}}(x^{0})$. Let $x\in S_{\alpha_{0}}(x^{0})$ and $\epsilon>0$, we want to show that there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$. Let $\beta_{0}\leq\eta<\alpha_{0}$ be such that $d(x_{\eta},y)<\epsilon/2$. By property (9) and Remark 5.2, it follows that there exists a strong $(\epsilon/2,d)$-chain $x=y_{0},y_{1},\ldots,y_{n-1},y_{n}=x_{\eta}$ from $x$ to $x_{\eta}$ whose points belong to $S_{\alpha_{0}}(x^{0})$. By the triangle inequality,